OpenStudy (abbycross167):

can someone please help me with this math question? please I'll fan and give a medal...

OpenStudy (abbycross167):

show the steps to graph this equation 6x < 24 + 3y

OpenStudy (mathmale):

There are more than just one way to graph this inequality (not equation, by the way). The intercept method is one of the fastest: Temporarily replace the < symbol with = Find the y-intercept: Let x=0, and solve the resulting equation for y. Write (x,y). Find the x-intercept: Let y=0, and solve the resulting equation for x. Write (x,y). Draw a dashed line thru these two intercepts. Choose a test point, e. g., (0,0). Substitute 0 for x and 0 for y. Is the resulting inequality true or false? If true, shade the side of your dashed line that includes the test point (0,0). Otherwise, shade the other side of your dashed line. If you'll show your work, I'll gladly give you further feedback.

OpenStudy (abbycross167):

@mathmale I have a question-- when you said to find the y-intercept and Let x=0, would I write the equation like this- X=0 6(0)=24+3y

OpenStudy (peachpi):

yes, then you'd solve that for y

OpenStudy (abbycross167):

so would y equal this? y>2x−8 ?

OpenStudy (abbycross167):

@peachpi ?

OpenStudy (peachpi):

yes, but that just makes the numbers smaller. You still have to follow @mathmale's steps

OpenStudy (mathmale):

If you set x=0 as I suggested, then you're left with 0 = 24 + 3y. Solve that for y. x should no longer appear in this equation; you've set it = to 0 already.

OpenStudy (mathmale):

If 0 = 24 + 3y, then 3y = -24. Find y. Then write (x,y), the y-intercept.

OpenStudy (abbycross167):

y=-8

OpenStudy (mathmale):

Yes. Your y-intercept is thus (0,-8). (because we purposely set x=0).

OpenStudy (mathmale):

Now find your x-intercept. Set y=0 and find x. Write (x,y).

OpenStudy (abbycross167):

ok so would it look like this? 6x=24+0

OpenStudy (mathmale):

Yes, and then x=4 when y=0. Write out the coordinates of this x-intercept, please.

OpenStudy (abbycross167):

(4,0)

OpenStudy (mathmale):

Yes Plot these 2 points. Draw a line thru them. Choose a test point (as I explained earlier) to determine which side of your dashed line you should shade. Can you finish this problem on your own with the above info?

OpenStudy (mathmale):

You will be finished when you've plotted these 2 points, drawn a dashed line through them, and then determined which side of the line should be shaded. Example:

OpenStudy (abbycross167):

Thank you soo much!! I have 3 more problems would you mind checking my steps?